find the length and width of arectangle
area=4square inches
perimeter=8 inches

Answer by edjones(7311)

wL=A
2w+2L=P
.
A) wL=4
B) 2w+2L=8
.
A) L=4/w
B)
2w+2(4/w)=8 Substitute
2w + 8/w = 8
2w^2+8=8w multiply each side by w to eliminate fractions.
2w^2-8w+8=0
2(w^2-4w+4)=0
(w-2)^2=0
w=2"
L=4/w=4/2=2"
.
Ed